OpenStudy (anonymous):

solve each equation for variable. 2log X= log 25

OpenStudy (jojo4eva):

x=5

OpenStudy (anonymous):

but how? :\

OpenStudy (anonymous):

thankss!!!

OpenStudy (jojo4eva):

go there and you will find it i have to go so click on that link and it will explain everything

OpenStudy (jojo4eva):

no probs

OpenStudy (jdoe0001):

\(\bf 2log(x)= log 25\qquad \textit{log rule of }a\ log(b)=log(b^a)\\ \quad \\ {\color{red}{ 2}}log(x)= log 25\implies log(x^{\color{red}{ 2}})= log(25)\implies log(x^{\color{red}{ 2}})= log(5^{\color{red}{ 2}})\)

OpenStudy (anonymous):

thank youu!!

OpenStudy (jdoe0001):

yw